OpenStudy (anonymous):

sine of pi divided by two times cosine of pi divided by seven plus cosine of pi divided by two times sine of pi divided by seven.

OpenStudy (anonymous):

\[\sin(\frac{\pi}{2})\times \cos(\frac{\pi}{7})\]?

OpenStudy (anonymous):

@satellite73 Write the expression as either the sine, cosine, or tangent of a single angle